F1-15 kN 40° F2- 26 kN 13 12 30° F3-36 kN

A.

Express F1 as a Cartesian vector.

B. Express F2 as a Cartesian vector.

C. Express F3 as a Cartesian vector.
Answer #1

(A) F1 = (15*sin40 i + 15*cos40 j) kN = (9.64 i + 11.50 j) kN

(B) and (C) solutions are in solved in sheet.

F, 26 KN づ 2 22.6 So 240 tooj) KN 3 = (31.21-18-5) KN. .92